Distant line Flectric dipole antenna LCoscillator 5434.35
Answer #1

wavelength is given by:

wavelength = c/f

c = speed of light = 3*10^8 m/sec

f = resonance frequency of RLC circuit

resonance frequency when

XL = Xc

w*L = 1/(w*C)

Since w = 2*pi*f

2*pi*f*L = 1/(2*pi*f*C)

f = 1/(2*pi*sqrt (LC))

f = 1/(2*pi*sqrt (0.294*10^-6*28.3*10^-12))

f = 55176385.68 Hz

Now wavelength of wave will be:

wavelength = 3*10^8/55176385.68

wavelength = 5.44 m

Please Upvote.

Comment below if you've any query.